OpenStudy (anonymous):

Which of the following is equivalent to the given expression? (Sin^2x)/(1-cosx) a. tanx + sinx b. cscx + cotx c. cotx sinx + tanx d.1 + cosx e. tanx cotx - cosx

OpenStudy (campbell_st):

start with the numerator \[\sin^2(x) + \cos^2(x) = 1~~~~~then~~~~~\sin^2(x) = 1 - \cos^2(x)\] which is the difference o 2 squares. ( (1 - cos(x))(1 + cos(x)) make the substitution and simplify